OpenStudy (anonymous):

a beam has a total volume of 60 ft cubed. find the max horizontal gap that the beam can span

OpenStudy (dumbcow):

V = L*W*H lets say L is horizontal length L = V/(W*H) L = 60/(W*H) inverse relationship, L increases as W and H decrease without any other restrictions, technically L goes to infinity as W or H goes to 0 so there is no max number, it can span any given length